# Fuzzy Prediction Sets: Conformal Prediction with E-values

### v2

@@ Abstract
We make three contributions to conformal prediction. First, we propose fuzzy conformal prediction sets that offer a degree of exclusion, generalizing beyond the binary inclusion/exclusion offered by classical prediction sets. We connect fuzzy prediction sets to e-values to show this degree of exclusion is equivalent to an exclusion at different confidence levels, capturing precisely what e-values bring to conformal prediction. We show that a fuzzy prediction set is a predictive distribution with an arguably more appropriate error guarantee. Second, we derive optimal conformal prediction sets by interpreting the minimization of the expected measure of a prediction set as an optimal testing problem against a particular alternative. We use this to characterize exactly in what sense traditional conformal prediction is optimal, and show how this may generally be used to construct optimal (fuzzy) prediction sets. Third, we generalize the inheritance of guarantees by subsequent minimax decisions from prediction sets to fuzzy prediction sets. All results generalize beyond the conformal setting to prediction sets for arbitrary models. In particular, we find that constructing a (fuzzy) prediction set for a model is equivalent to constructing a test (e-value) for that model as a hypothesis.
Keywords: Conformal prediction, prediction set, e-value, fuzzy confidence set, certified decision.
@@ 1 Introduction
1.1 Traditional conformal prediction
@@
To construct such a prediction set, the idea behind conformal prediction is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the prediction set is formed by collecting the values $z$ for which the hypothesis is not rejected at $\alpha$:
@@
Figure 1: A traditional and fuzzy conformal prediction set. The solid line represents a traditional level $\alpha=0.05$ prediction set: the points $z_{n+1}$ at which this line is zero form the prediction set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued prediction set, offering a degree of exclusion at each point.
1.2 Fuzzy conformal prediction
A downside of a traditional conformal prediction set is that it is binary in nature: a sample point $z$ either falls inside the set or not. To tackle this, we propose fuzzy conformal prediction sets.
For each point $z\in\mathcal{Z}$, a traditional prediction set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion-exclusion to a degree of exclusion, in the interval $[0,1]$. We illustrate this in Figure 1, where the solid line represents a traditional prediction set and the dashed line a fuzzy prediction set.
@@ 1.2 Fuzzy conformal prediction
For a binary test, this collapses to the classical coverage guarantee (1), since $C_{\alpha}^{Z^{n}}(z_{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{Z^{n}}\}$.
1.3 E-values and fuzzy prediction
An apparent downside of a fuzzy prediction set is that the ‘degree of exclusion of $z$ at level $\alpha$’ is difficult to interpret. We overcome this by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different significance levels. This relies on the recently established equivalence between fuzzy tests and ‘e-values’ (koning2025continuoustestingunifyingtests).
To facilitate this, we rescale the prediction set by dividing it by $\alpha$:
@@ 1.3 E-values and fuzzy prediction
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy prediction set $\mathcal{E}^{Z^{n}}(z)$ at any point $z$ may be directly interpreted as (the reciprocal of) the smallest significance level at which $z$ is excluded, under a generalized Type I error guarantee (grunwald2022beyond; koning2024post). This replaces the ‘degree of exclusion of a point $z$ at level $\alpha$’ by an ‘exclusion at level $\alpha^{\prime}=1/\mathcal{E}^{Z^{n}}(z)$’. Because of the rescaling, the guarantee (2) must be reformulated to
$\mathrm{E}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1,$
which also enables $0,\infty]-$valued fuzzy prediction sets.
We illustrate fuzzy prediction sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1 that also includes a $0,\infty]-$valued fuzzy prediction set. Here, we see that the traditional prediction set only offers an exclusion at level $1/20=0.05$, whereas the fuzzy prediction sets offer exclusions at various significance levels for different points in the sample space. Compared to the traditional prediction set, the unbounded fuzzy prediction set expresses very strong evidence against large values of $z_{n+1}$. The bounded version is capped at evidence corresponding to a rejection at level $1/20=0.05$, but trades this against more evidence for small values of $z_{n+1}$ compared to the unbounded version.
We show that fuzzy prediction sets may easily be merged, by averaging their pointwise evidence under arbitrary dependence, or multiplying under independence. Applying this averaging to non-fuzzy prediction sets generally produces fuzzy prediction sets.
Figure 2: One traditional and two fuzzy prediction sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ prediction set, the dashed line represents a fuzzy $0,1/\alpha]-$valued prediction set, and the dotted line a fuzzy $0,\infty]-$valued prediction set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=a$ corresponds to an exclusion at significance level $1/a$.
@@ 1.4 Optimal (fuzzy) prediction sets
Besides generalizing to fuzzy prediction sets, we also contribute to classical non-fuzzy conformal prediction by studying its optimality. In fact, our optimality results generalize beyond exchangeability to prediction sets for arbitrary models.
@@
The key to unlocking our optimality results is to show that a prediction set for a model $\mathcal{P}$’ is equivalent to a single test for the hypothesis $\mathcal{P}$. We subsequently link optimal prediction sets to optimal testing theory. In particular, we show that a prediction set is of minimal expected size under some measure $\mu$,
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{Z^{n}}[\mu(C_{\alpha})],$
if and only if the corresponding test is most powerful against some alternative hypothesis that incorporates $\mu$.
In Section 5, we specialize the optimality to the classical conformal prediction setting where $\mathcal{P}$ is the exchangeable model to derive optimal prediction sets for conformal prediction. We use this to show in what sense classical conformal prediction is optimal, given a choice of conformity score.
@@
These results generalize to fuzzy prediction sets, which are equivalent to e-values. We connect optimal fuzzy prediction sets to expected-utility optimal e-values koning2025continuoustestingunifyingtests. Classical conformal prediction is recovered as a special case, for the ‘Neyman–Pearson utility function’ $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$, when the e-value corresponds to the Neyman–Pearson most powerful test at level $\alpha$. Other utility functions may be used to express different preferences regarding the value of different amounts of evidence, which subsequently shapes the fuzzy prediction set.
1.5 Certifying subsequent decisions
In Section 6, we explain how a fuzzy prediction set can be used in subsequent decision making. Here, we show how the coverage guarantee offered by a fuzzy prediction set may be inherited by subsequent minimax decisions. This generalizes ideas of kiyani2025decision from traditional conformal prediction sets to fuzzy conformal prediction sets and general prediction sets.
@@ 1.5 Certifying subsequent decisions
Most of our findings here build on and extend results of grunwald2023posterior and andrews2025certified to conformal prediction. This relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior; an object recently studied by grunwald2023posterior. This shows that the ‘E-posterior’ is deeply connected to classical hypothesis testing. Furthermore, our perspective unifies the notions of P-certification and E-certification that were recently developed by andrews2025certified.
1.6 Covariates and application
In Section 7, we show how our results extend to the setting of covariates. In the covariate setting, the tuple $Z^{n+1}$ can be decomposed into $Z^{n+1}=(X^{n+1},Y^{n+1})$, and we measure the covariate $X_{n+1}$ alongside $Z^{n}$. The goal is then to obtain a prediction set for $Y_{n+1}$. Although the use of covariates is standard in conformal prediction, we defer their inclusion to this later section as they add considerable notational overhead and are not essential to our main results.
@@ 1.6 Covariates and application
We apply our methodology to a classical character-recognition problem in Section 8. Here, we showcase what fuzzy prediction sets offer over classical prediction sets, and study how different utility functions shape the fuzzy prediction sets.
2 Related literature
@@
To the best of our knowledge, fuzzy confidence sets were first explicitly proposed by geyer2005fuzzy for confidence sets on parameter spaces. Compared to their work, we deliver several innovations. First, while conceptually the same type of object, we stress that our work is on fuzzy prediction sets, which may be viewed as confidence sets on a sample space instead of on a parameter space. Second, geyer2005fuzzy present fuzzy confidence sets as a solution to under-coverage of classical Neyman–Pearson optimal confidence sets for discrete data in small samples. We advocate for a much wider use, beyond discrete data, by using a generalization of the Neyman–Pearson framework from the recent e-value literature. Third, we show that a fuzzy ‘degree of exclusion at level $\alpha$’ is equivalent to an exclusion at a different significance level under a generalized Type I error guarantee. We believe this may make fuzzy prediction/confidence sets more palatable, as this makes them easier to interpret and work with. Finally, we also attempt to show the relevance of fuzzy prediction sets by showing how they may be used in subsequent decision making, further motivating their relevance over non-fuzzy prediction sets.
@@
As the literature on conformal prediction and e-values remains fairly small, we attempt to give a complete overview. According to vovk2024conformal, precursors of conformal prediction (Gammerman1998; Vovk1999) actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that prediction sets were more naturally obtained by using p-values. This is in line with our finding that e-values are equivalent to fuzzy prediction sets, which implies that e-values play no fundamental role in classical (non-fuzzy) prediction sets.
@@
The first explicit connection between modern conformal prediction and e-values appears in the 2020 arXiv version of vovk2024conformal, where he exploits the merging properties of e-values to enable cross-conformal prediction by averaging over data splits. Since then, several other works have used e-values as a pragmatic tool to construct conformal prediction sets. For example, in the context of conformal selection, lee2024boosting, lee2025full, lee2025selection and nair2025diversifying compute e-values for (non-)selection events and plug these into the e-BH procedure to control the false discovery rate. gauthier2025values averages conformal e-values over ambiguous ground truths, where observations may carry multiple valid labels simultaneously. Moreover, gauthier2025values also considers e-values to select a prediction set with a fixed number of elements. Our work differs from this line of work, as our intention is to directly report the e-values as output in the form of a fuzzy prediction set, rather than using them as an intermediate step to obtain classical prediction sets.
@@
In the early version, vovk2024conformal was also the first to propose using conformal e-values of the form
@@
where $T$ is a non-negative statistic. This form was later independently rediscovered in the first arXiv version of wang2022false, koning2023post and balinsky2024enhancing, and proven to be the only admissible type of e-values for conformal prediction by koning2023post. A key issue of the form (3) is that there is little guidance on how to appropriately select this statistic $T$ and gauthier2025values find the outcomes to be highly sensitive to this choice. We provide guidance by introducing a general expected-utility framework to conformal prediction, which enables users to quantify their preference for different degrees of evidence, and allows them to specify a measure $\mu$ that determines the most important regions of the outcome space. This nests classical conformal prediction as a special case for a particular choice of the utility function, as well as the setting studied by vovk2024conformal for a logarithmic choice of utility function and counting measure $\mu$ in classification problems. In our application, we illustrate that log-utility is not always appropriate in the conformal prediction context. This is not surprising, as log-utility is generally considered in long-run sequential data-collection settings instead of the single-batch setting we consider here.
3 Prediction sets
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. [^1] This is easily generalized to more general joint sample spaces, but we stick to copies of the same sample space for the sake of exposition. A prediction set is classically defined as a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
@@ 3 Prediction sets
Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability in Section 5. Moreover, we postpone the inclusion of covariates to Section 7 as their inclusion considerably clutters the notation and is not relevant for the discussion here.
@@ 3.1 Slicing a prediction set for $Z^{n+1}$
We illustrate this in Figure 3, where union of the slices $\{C_{\alpha}^{z^{n}}\}_{z^{n}\in\mathcal{Z}^{n}}$ forms the prediction set $C_{\alpha}$. Here, we stress that the prediction set $C_{\alpha}$ is deterministic: the randomness in $C_{\alpha}^{Z^{n}}$ is purely due to the random realization of the point $Z^{n}$ at which it is ‘sliced’.
The validity of $C_{\alpha}$ as a prediction set for $Z^{n+1}$ under the model $\mathcal{P}$ is equivalent to the validity of $C_{\alpha}^{Z^{n}}$ in the sense of (4). Indeed, by (5) we have $\mathbb{I}\{Z^{n+1}\in C_{\alpha}\}=\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}$, so that
@@
for any $P$. As a consequence, we may focus our attention to constructing prediction sets $C_{\alpha}$ for $Z^{n+1}$.
@@
These arguments may be extended to slicing based on any part or statistic of $Z^{n+1}$ that is observed. We illustrate this in the covariates setting described in Section 7 where $Z_{n+1}=(X_{n+1},Y_{n+1})$ and we observe $X_{n+1}$ alongside $Z^{n}$, to construct a prediction set for $Y_{n+1}$. There, we still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$ but then slice it at $(Z^{n},X_{n+1})$.

### v3

@@ Abstract
Prediction sets offer a binary inclusion/exclusion for each element at the same fixed confidence level. We generalize to fuzzy prediction sets, which exclude elements at their own data-driven confidence level. Our key insight is that a fuzzy prediction set is an e-value, capturing precisely what e-values bring to predictive inference. Fuzzy prediction sets inherit the merging properties of their e-value, offer richer guarantees to decision-makers. We also show in what sense optimal e-values give rise to optimal (fuzzy) prediction sets. We apply our results to conformal prediction, deriving optimal fuzzy conformal prediction sets, and characterizing in what sense classical conformal prediction is optimal.
@@ 1 Introduction
1.1 Traditional prediction sets
Prediction sets, with conformal prediction sets at the front, have been one of the main success stories in bringing uncertainty quantification to machine learning predictions.
@@ 1.1 Traditional prediction sets
To construct such a prediction set, the idea is to plug-in hypothetical values $z$ for $Z_{n+1}$ and test whether $(Z_{1},\dots,Z_{n},z)$ is exchangeable at level $\alpha$. Repeating this exercise for each plug-in value $z$ in the sample space, the prediction set is formed by collecting the values $z$ for which the hypothesis is not rejected at level $\alpha$:
@@
Figure 1: A traditional and fuzzy prediction set. The solid line represents a traditional level $\alpha=0.05$ prediction set: the points $z_{n+1}$ at which this line is zero form the prediction set, and the points at which it equals 11 its complement. The dashed line represents a fuzzy $[0,1]$-valued prediction set, offering a degree of exclusion at each point.
1.2 Fuzzy prediction sets
A downside is that a prediction set is binary: a sample point $z$ either falls inside the set or not. We overcome this by generalizing to fuzzy prediction sets.
A traditional prediction set may be reinterpreted as a function $C_{\alpha}^{Z^{n}}:\mathcal{Z}\to\{0,1\}$, which outputs an inclusion (0) or exclusion (1) for each point $z\in\mathcal{Z}$. A fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}:\mathcal{Z}\to[0,1]$ generalizes beyond the binary inclusion/exclusion to a degree of exclusion. We illustrate this in Figure 1, where the solid line represents a traditional prediction set and the dashed line a fuzzy prediction set.
@@ 1.2 Fuzzy prediction sets
For a binary test, this collapses to the classical guarantee (1), as $C_{\alpha}^{Z^{n}}(z_{n+1})=\mathbb{I}\{z_{n+1}\not\in C_{\alpha}^{Z^{n}}\}$.
1.3 E-values and fuzzy prediction sets
We connect fuzzy prediction sets to e-values (Grünwald et al., 2024; Vovk and Wang, 2021; Howard et al., 2021; Shafer, 2021; Koning, 2025a) by merging the degree of exclusion into the significance level, allowing different points $z$ to be excluded at different levels.
To facilitate this, we rescale from $[0,1]$ to $0,1/\alpha]$ by dividing the prediction set by $\alpha$:
@@ 1.3 E-values and fuzzy prediction sets
Due to the rescaling, $\mathcal{E}^{Z^{n}}(z)=1/\alpha$ now corresponds to a full exclusion at level $\alpha$. More generally, we show that the value of a fuzzy prediction set $\mathcal{E}^{Z^{n}}(z)$ at any point $z$ may be directly interpreted as an exclusion at the data-driven level $1/\mathcal{E}^{Z^{n}}(z)$ under an extension of the Type I error to data-dependent levels (Grünwald, 2024; Koning, 2024). Because of the rescaling, the guarantee (2) must be reformulated to $\mathrm{E}[\mathcal{E}^{Z^{n}}(Z_{n+1})]\leq 1$, which also enables $0,\infty]-$valued fuzzy prediction sets.
We illustrate fuzzy prediction sets on this evidence scale in Figure 2, which is a rescaled version of Figure 1 that also includes a $0,\infty]-$valued fuzzy prediction set. Here, the traditional prediction set only offers exclusions at level $1/20=0.05$, whereas the fuzzy prediction sets offer exclusions at various significance levels for different points $z$ in the sample space.
The key insight that underlies this is that a fuzzy prediction set $\mathcal{E}^{Z^{n}}$, is equivalent to an e-value $\mathcal{E}(Z^{n},Z_{n+1}):=\mathcal{E}^{Z^{n}}(Z_{n+1})$. Moreover, we show that the fuzzy prediction set $\mathcal{E}^{Z^{n}}$ is valid for a model $\mathcal{P}$ if and only if this e-value $\mathcal{E}$ is valid for the hypothesis $\mathcal{P}$. This specializes to classical prediction sets and tests, which are $\{0,1/\alpha\}$-valued e-values.
We use the equivalence to e-values to show that fuzzy prediction sets inherit their flexible merging properties: they may be averaged under arbitrary dependence, and multiplied under independence. Averaging non-fuzzy prediction sets generally produces fuzzy prediction sets. Multiplying prediction sets yields their union, and the corresponding confidence level is the product of the individual levels.
Figure 2: One traditional and two fuzzy prediction sets on the evidence scale. The solid line represents a traditional level $\alpha=0.05$ prediction set, the dashed line represents a fuzzy $0,1/\alpha]-$valued prediction set, and the dotted line a fuzzy $0,\infty]-$valued prediction set. An evidence value of $\mathcal{E}^{Z^{n}}(z)=e$ corresponds to an exclusion at the data-driven confidence level $1/e$.
@@ 1.4 Optimal (fuzzy) prediction sets
The equivalence to an e-value allows us to connect optimal prediction sets to the rich statistical theory on optimal hypothesis testing. In particular, we show that a prediction set is of minimal expected size under some measure $\mu$,
$\operatorname*{arg\,min}_{C_{\alpha}}\mathrm{E}^{Z^{n}}[\mu(C_{\alpha}^{Z^{n}})],$
@@
if and only if the corresponding test is most powerful against an alternative that involves $\mu$. This most powerful test may be viewed as an e-value that optimizes the expected ‘Neyman–Pearson utility function’ $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$ (Koning, 2025a). Other utility functions may be used to express different preferences regarding the value of different amounts of evidence.
We specialize this to conformal prediction in Section 4, leveraging recent results on optimal e-values for exchangeability (Koning, 2025b). This also reveals in what sense classical conformal prediction is optimal, given a choice of conformity score.
1.5 Decision making, application and covariates
@@
To motivate fuzzy prediction sets, we show that they offer richer loss bounds to decision-makers than classical prediction sets. For this purpose, we generalize the framework of Kiyani et al. (2025) for traditional conformal prediction sets. One of our approaches connects to Grünwald (2023), and relies on the surprising (to us) observation that a fuzzy confidence set is equivalent to (the reciprocal of) an E-posterior. This shows that the E-posterior is deeply connected to classical hypothesis testing. Furthermore, we unify the related notions of P-certification and E-certification that were recently developed by Andrews and Chen (2025).
@@
We apply our methodology to character-recognition in Section 6. Here, we showcase the richness of classical prediction sets, and study how different utility functions shape the fuzzy prediction sets. In Appendix E, we extend to covariates.
1.6 Related literature
@@
To the best of our knowledge, fuzzy confidence sets were first proposed by Geyer and Meeden (2005) as a solution to under-coverage of classical Neyman–Pearson optimal confidence sets for discrete data in small samples. We instead advocate for a much wider use, by using a generalization of the Neyman–Pearson framework through e-values. A key innovation compared to Geyer and Meeden (2005) is that a fuzzy ‘degree of exclusion at level $\alpha$’ is equivalent to an exclusion at a data-driven significance level, which we believe makes fuzzy prediction/confidence sets much more palatable.
@@
As the literature on conformal prediction and e-values remains fairly small, we attempt to give a complete overview. According to Vovk (2025), precursors of conformal prediction (Gammerman et al., 1998; Vovk et al., 1999) actually relied on e-values (under a different name). He attributes their demise in conformal prediction to the fact that prediction sets were more naturally obtained by using p-values. This is in line with our finding that e-values are equivalent to fuzzy prediction sets, which implies that e-values play no fundamental role in classical (non-fuzzy) prediction sets.
@@
The first explicit connection between modern conformal prediction and e-values appears in the 2020 arXiv version of Vovk (2025), where he exploits the merging properties of e-values to enable cross-conformal prediction by averaging over data splits. Since then, several other works have used e-values as a pragmatic tool to construct conformal prediction sets. For example, in the context of conformal selection, Lee and Ren (2024), Lee et al. (2025), Lee and Ren (2025) and Nair et al. (2025) compute e-values for (non-)selection events and plug these into the e-BH procedure to control the false discovery rate. Gauthier et al. (2025) average conformal e-values over ambiguous ground truths, where observations may carry multiple valid labels simultaneously. Moreover, Gauthier et al. (2025) also consider e-values to select a prediction set with a fixed number of elements. Our work differs from this line of work, as our intention is to directly report the e-values as output in the form of a fuzzy prediction set, rather than using them as an intermediate step to obtain classical prediction sets.
@@
In the early version, Vovk (2025) was also first to propose conformal e-values of the form
@@
where $T$ is a non-negative statistic. This form was later independently rediscovered in the first arXiv version of Wang and Ramdas (2022), Koning (2025b) and Balinsky and Balinsky (2024), and proven to be the only admissible types of e-values for conformal prediction by Koning (2025b).
A key issue of the form (3) is that there is little guidance on how to appropriately select the statistic $T$ and Gauthier et al. (2025) find the outcomes to be highly sensitive to this choice. We turn away from (3), and instead introduce a general expected-utility optimality framework to conformal prediction, which enables users to quantify their preference for different degrees of evidence, and allows them to specify a measure $\mu$ that determines the most important regions of the outcome space. This can be interpreted as a way to optimally select $T$ in (3). Our framework nests classical conformal prediction as a special case for a particular choice of the utility function, as well as the setting studied by Vovk (2025) for a logarithmic choice of utility function and counting measure $\mu$ in classification problems. In our application, we illustrate that log-utility is not always appropriate in the conformal prediction context. This is not surprising, as log-utility is generally considered in long-run sequential data settings instead of the single-batch setting common in conformal prediction.
2 Relationship between prediction sets and tests
2.1 Prediction sets
Let $\mathcal{Z}$ denote the sample space of a single observation, and $\mathcal{Z}^{n+1}$ the joint sample space of $(n+1)$ observations. [^1] This is easily generalized to more general joint sample spaces, but we stick to copies of the same sample space for the sake of exposition. A prediction set is a set $C_{\alpha}^{Z^{n}}$ with the level $\alpha>0$ coverage guarantee
@@ 2.1 Prediction sets
Conformal prediction is the special case where the model $\mathcal{P}$ is the collection $\mathcal{P}^{\textnormal{exch.}}$ of exchangeable distributions on $\mathcal{Z}^{n+1}$, but we delay specializing to exchangeability to Section 4. Moreover, we postpone the inclusion of covariates to Appendix E as their inclusion considerably clutters the notation and is not relevant for the discussion here.
@@ 2.2 Slicing a prediction set for $Z^{n+1}$
We illustrate this in Figure 3, where the union of the slices $\{C_{\alpha}^{z^{n}}\}_{z^{n}\in\mathcal{Z}^{n}}$ forms the prediction set $C_{\alpha}$. Here, we stress that the prediction set $C_{\alpha}$ is deterministic: the randomness in $C_{\alpha}^{Z^{n}}$ is purely due to the random realization of the point $Z^{n}$ at which it is ‘sliced’.
Validity of $C_{\alpha}$ as a prediction set for $Z^{n+1}$ under the model $\mathcal{P}$ is equivalent to the validity of $C_{\alpha}^{Z^{n}}$ in the sense of (4). Indeed, by (5) we have $\mathbb{I}\{Z^{n+1}\in C_{\alpha}\}=\mathbb{I}\{Z_{n+1}\in C_{\alpha}^{Z^{n}}\}$, and so
@@
for any $P$. As a consequence, we may focus on constructing prediction sets $C_{\alpha}$ for $Z^{n+1}$.
@@
These arguments may be extended to slicing based on any part or statistic of $Z^{n+1}$ that is observed. We illustrate this in the covariates setting described in Appendix E where $Z_{n+1}=(X_{n+1},Y_{n+1})$ and we observe $X_{n+1}$ alongside $Z^{n}$, to construct a prediction set for $Y_{n+1}$. There, we still construct a prediction set $C_{\alpha}$ for $Z^{n+1}$ but then slice it at $(Z^{n},X_{n+1})$.